OpenStudy (anonymous):

Given the following two points, (a) find the distance between them, (b) find the midpoint, and (c) submit the graph to the Dropbox. (-4, 1), (0, -3)

OpenStudy (anonymous):

Distance: \[\sqrt{(-4-0)^2 + (1 - (- 3))^2} = \sqrt{16+16} = 4\sqrt{2}\] Midpoint: \[(\frac{-4+0}2, \frac{1-3} 2) = (-2, -1)\]